two identical charges having the charge 2*10^(-4) C,and having the mass 10g are kept at separation by 10cm.what would be the velocity of particles when they the separation is maximum.

Grade:12

1 Answers

vikas askiitian expert
13 years ago

initial potential energy is (PE)i = kq2 /r                  (charges are of same magnitude q)

                    k = 9 . 109 units

                                        (PE)i  =3600j                 (after substituting value of charge and K)

finally when distance is maximum then (PE)=kq2 /r                    

                                r=r(max)=infinity

      so final potential energy is 0...

initially kinetic energy is zero ....

finally kinetic energy is (KE)f =  mv12 /2   +  mv2 2 /2

                                         =m(v12 + v22 )/2 ................1

from conservation of energy ,total energy initial = total energy final

    therefore ,     m(v12 +v22 )/2 =3600 ..................2

       since no external; force is acting so momentam will be conserved..

     initial momentam =0

    final momentam=mv1 +mv2

      mv1 +mv2 =0

            v1=-v2 ..................1

    now from eq 2 putting v1=-v2

     v1=600m/s

      v2=-600m/s